

********************************************************************
UARS WEEKLY STATUS REPORT - 29 JAN 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
data without apparent problems.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Jan 26 and is scheduled to operate though
Feb 11. The MLS team is operating the instrument in a mode emphasizing
scientific observations at the northern latitudes.
***********
HRDI
***********
HRDI continues to operate normally.
***********
WINDII
***********
WINDII is currently on, and operating normally. It is expected to
continue measurements until February 7.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
0 days HAL (4.80) PROD (L3) (1874R-1721C-1F=152L)
147 days SUS (7.00) PROD (L3) (2051R-1739C-9F=303L)
3 days HAL (4.81) ENG (L1-L3) (3R-3C=0L)
0 days WIN (5.11) ENG (L1-L3) (1R-1F=1L)
Upgrades to UCSS10.1, UOAS5.4, MSS4.1 and NRT 2.6 are completed. A message
pertaining to logical changes required by RACs using "nrt watch"
programs was sent to UARS distribution.
CDHF completed the upgrade of LSE to Version 4.5 and PCA to
version 4.5-4 as part of the ongoing Y2K implementation effort.
SOLSTICE Wall Clock time increase.- Enough debug output has been produced to
enable tracking down the problem on the operational system where job step 2
takes an inordinate amount of time to complete. Steps are underway to create an
engineering job for this purpose.
DAAC data created for HALOE, HRDI, NMC, PEM, and UKMO in the month
of November 1998 has been loaded into the DAAC Transfer Tables.
CDHF pending software upgrades: IDL 5.03(installed/not operational),
OpenIngres, HALOE 4.81.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 January 1999 (Orbit 40152, GMT Day 016) through 22
January 1999 (Orbit 40255, GMT Day 022).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 022/2052 40254 PEM Test
The beta angle ranged between +14.6 degrees to 43.3 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 74.0
degrees on 31 January 1999 (DOY 011).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.7 to 1.3 degrees C. Additional battery performance data
for each battery on 15 January 1999 and 22 January 1999 are provided in
Appendix D.
The clock error ranged from -10.1 to 11.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 01
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% PEM 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/362 - 2287.496912
Transponder B: 99/357 - 2287.499093
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/021 show the observatory data loss to be
21 hours, 50 minutes, 20.788 seconds (an increase of 16.480 seconds since
last report period). This is a 0.0340 percent data loss which equals a
99.9660 percent data capture for the mission.
The increase of 16.480 seconds occurred over four (4) days:
7 2.048 seconds on Day 015 due to hit in data (IP),
7 9.696 seconds on Day 016 due to insufficient overlap,
7 1.120 seconds on Day 018 due to hit in data (IP), and
7 3.616 seconds on Day 019 due to insufficient overlap.
The UARS Test and Training Simulator (UTTS) remains non-operational. The
estimated time to repair the UTTS and return it to operation is unknown.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
UARS RE-ENGINEERING - TELEMETRY PROCESSING - No change
MEETINGS
The FOT attended a UARS Y2K end-to-end testing meeting on 20 January.
Representatives of the various UARS related facilities (i.e. MSOCC, CDHF,
IPD, etc) were in attendance. As of this date, there are several entities
that are not quite ready to participate in UARS Y2K end-to-end testing.
One major show stopper for UARS is the current system disk problem with the
UTTS.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 20
February 1999.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
none
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 January 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
40166 016/2254 99-027 OBC Clk Adj fr 20.6 to 0.00 msec/day
40216 020/0745 99-028 OBC Clk Adj fr 0.00 to 20.6 msec/day
40243 022/0158 99-029 PEM HEP1 EPR ENA
40247 022/0940 99-030 OBC Clk Adj fr 20.6 to 0.00 msec/day
40254 022/2052 99-031 PEM Genascii
40254 022/2052 99-032 PEM PMON Action 26
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 January 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 40205,15-17
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 January 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
40166 016/22:59 (16 Jan 99) from 20.6 to 00.0 msec/day
40216 020/07:53 (20 Jan 99) from 00.0 to 20.6 msec/day
40247 022/09:43 (22 Jan 99) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 January 1999
15 Jan 1999 - GMT Day 015 Beta = 14.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +44.8/+22.4 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +112.0/+11.2 +7.09/+5.15 +32.0/-15.2 23.8 1.023 23.1
3 +22.4/-22.4 +5.47/+3.87 +31.6/-16.8 23.7 1.022 23.0
22 Jan 1998 - GMT Day 022 Beta = 43.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +39.2/+11.2 -3.05/-4.31 +0.0/ +0.0 16.8 0.000 0.0
2 +22.4/-50.4 +6.44/+4.51 +29.6/-16.0 25.1 1.023 22.1
3 +33.6/-11.2 +5.15/+3.87 +28.8/-16.0 25.1 1.026 21.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 22 January 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-001 FOT see list Jan 98 14170 Generic Late Acq
SEVNTFW99/04